OpenStudy (anonymous):

The distance formula can be used to prove a quadrilateral has a right angle congruent angles parallel sides congruent sides

hartnn (hartnn):

distance formula gives you distance between 2 points. which means it gives you length of the side of a quadrilateral. using distance formula, you can prove that 2 sides equal.

OpenStudy (anonymous):

Thank you so much (; CAn you help me with a few more? @hartnn

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

Line AB contains points A (3, −2) and B (1, 8). The slope of line AB is −5 negative 1 over 5 1 over 5 5

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

hartnn (hartnn):

here x1 = 3 y1 =-2 x2= 1 y2 =8

OpenStudy (anonymous):

@hartnn Is it B?

OpenStudy (anonymous):

it has to be negative right

hartnn (hartnn):

nopes, what you got for numerator ? y1-y2 = -2-8 = ... ?

OpenStudy (anonymous):

-5 @hartnn

hartnn (hartnn):

\(\Large \checkmark \)

OpenStudy (anonymous):

@hartnn your the best (: !! I have quite a few more if you want to help me?

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

Find the x value for point I such that DI and EI form a 3:5 ratio. Segment DE is shown. D is at negative 3, 2. E is at 5, 5. 0 3 −1 1

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):